# Benchmarking Misuse Mitigation Against Covert Adversaries

### v1

@@ Appendix A Additional related work
Dangerous capability evaluations. Dangerous capability evaluations attempt to estimate the proficiency of frontier models on tasks where language models could unlock large scale harm, for example, cyber-offense, persuasion, bio-engineering, and self-replication [24, 50, 51]. Frontier model developers most often conduct dangerous capability evaluations internally and report high-level results via system cards [52, 53, 54, 47]. Dangerous capability evaluations are run under a threat model where the human attempting misuse is either directly querying the model (typically with safeguards like safety training removed) or applying an undisclosed jailbreak or elicitation method. Sometimes dangerous capability evaluations are paired with human uplift studies, which evaluate the extent that a language model helps humans perform dangerous or dual-use tasks [55, 56]. In contrast, our threat model assumes that model developers will deploy standard safeguards and that attackers will attempt to subvert safeguards via attack strategies like decomposition attacks and jailbreaking.
@@
Jailbreaking methods. Most jailbreaks try to coerce a model into eliciting disallowed content, e.g., “Tell me how to build a bomb”[28, 29, 25, 2, 30]. Many optimize for a fixed target string (“Here is how to build a bomb…”) [29, 25] and others look for non-refusal answers [2, 31, 26]. These approaches are usually benchmarked on questions whose answers are easy to find via the web [4, 5]. Outputs from jailbreaks, even when “successful,” often return vague or erroneous instructions [11]. HarmBench’s harder context-based tasks represent an attempt to alleviate this, yet are largely saturated by open-weight LLMs [4, 21]. Here, we instead measure misuse-uplift on genuinely hard, refused tasks and introduce BSD, which pairs uplift with an explicit detectability axis that is missing from refusal-only metrics. Similar to [10, 12], we show that jailbreaking prompts are relatively easy to detect, whereas decomposition attacks are significanttly harder to detect.
@@
Decompostion methods. Decomposition attacks, introduced in previous work [18, 19], are methods that use benign-looking sub-queries to help solve a malicious task. That said, [18] run a decomposition attack on a set of Python scripts generated by Claude 3 Opus and judged by GPT-4. We note that the provided example tasks are not refused by strong models, e.g. Claude Sonnet 3.5 or GPT-4o, and thus cannot be used to evaluate our misuse uplift threat model. Similarly, [18] does not compare decomposition attacks with established jailbreak methods. [19] studies the increase in their introduced Impermissible Information Leakage on WMDP, but as shown in Section 5, strong models directly answer these queries and decomposition harms accuracy, making WMDP a poor misuse proxy. By contrast, our study (i) frames decomposition as a way to evade detectability (Section 2), (ii) benchmarks the methods on a misuse-uplift metric that factors in both task difficulty and strong model refusal, and (iii) introduces improved decompositions that outperform prior work (Section 5).
@@
Stateful defenses. A parallel line of work shifts from single-prompt screening to sequence-level scrutiny. In computer vision, Stateful Detection compares each new input to a sliding window of earlier queries [32]; Blacklight speeds this up with locality-sensitive hashing [33], and PIHA swaps raw pixels for perceptual hashes to cut false positives [34]; and Mind-the-Gap augments the windowed distance test with adaptive thresholds yet still falls to the OARS adaptive attack [35, 36]. PRADA detects model stealing by flagging query sequences whose distances deviated from benign traffic [57]. Outside of vision, Clio clusters millions of conversation snippets to surface coordinated abuse, but publishes no quantitative evaluations and does not consider user-level defenses [37]. Our work (Section 4) proposes a detector for misuse uplift that uses a buffer to keep track fo the most concerning queries, and shows that even with mantaining a memory across many independent queries, decomposition attacks are harder to flag than standard jailbreaks.
@@ Appendix B Threat model details
    Open-weight models are currently weaker than proprietary models. Open-weight models— models with downloadable weights—have historically trailed proprietary systems in benchmark performance by at least 6 months [58]. While this performance gap is closing, it likely still holds for current frontier open-weight and closed-weight models [59, 60, 61].
@@
    Open-weight models can be made unsafe. The safety-training and guardrails on open-weights models can be removed with only modest additional fine-tuning [62, 63, 41]. While there is early work attempting to make models robust to fine-tuning attacks [64, 65], this problem is difficult— e.g., defense here is strictly harder than that for adversarial examples or jailbreaks [66].
@@ B.1 Alternative assumptions
Language model uplift is a proxy for human uplift. First, we note that helpful-only (unsafe) models may serve as cheap (but imperfect) substitutes for non-expert humans in a misuse evaluation. This means that our evaluations can provide information on human uplift [67]. [^3] We note that this is similar to the assumptions made in scalable oversight [68]. For example, a weaker model might serve as an imperfect stand-in for a human with beginning-to-intermediate software engineering ability [69] in a cyber-misuse setting. In this case, the helpful-only (unsafe) model would approximate a steps performed by a human attacker: reconnaissance and vulnerability discovery, weaponization, exploitation, escalation, etc. [70], delegating to the proprietary (safe) model when needed.
@@
Imagine the adversary can only run the model on a single CPU at roughly 1 token per second, whereas the defender hosts the same weights on a GPU that runs at 100 tokens per second. Jailbreaking the defender’s endpoint grants the attacker two orders of magnitude more effective compute per wall-clock hour. For agent and reasoning workflows where the model plans, branches, etc, this translates into substantially deeper search, which in turn has been shown to raise success rates on reasoning-intensive tasks [54].
@@
Unsafe stronger models can be complementary with safe weak models. Even in a world where the strongest models are willing to do harmful actions, the capabilities of these models may be complementary with those of proprietary models with safety training [64]. For example, while a helpful-only model may have vastly more world knowledge, it may still use a (weaker) safe proprietary model that has longer/more consistent reasoning to do more harm in an agent setting.
@@ Appendix C Cybersecurity dataset and results
We also provide cybersecurity misuse uplift results in Figure 7 with Qwen-2.5 7B as the attacking model for decomposition. The misuse uplift, how much the strong model augments the weaker model, is generally more modest in the cybersecurity setting compared to the uplift measured in the BSD biology dataset (see Figure 1). This due in-part because the baseline rate of refusal for cybersecurity questions are much lower across all models (see the directy query accuracy results in Figure 7).
Figure 7: The misuse rate for decomposition attacks vs direct querying for BSD cyber questions. While we find modest misuse uplift (the misuse rate increases over the dashed line), the attacker gains for BSD cyber questions are less than that for BSD bio questions (Figure 1)
@@ Appendix D Protocols for decomposition
We provide a high-level description of the different decomposition attack strategies we consider in this work in Table 2.
[…]
@@
[…]
@@ Appendix E BSD Pipeline Details
We generate BSD in three stages in order to get broad topic coverage. First, we generate 3000 candidate biology questions with GPT-4.1 [38] using the BSD Biology dataset [39] as seed questions. After filtering, this yields 15 total questions that meet our criteria of difficulty and harmfulness. Next, we identify 10 plausibly harmful topics in biology that are not covered in the existing question set. For the next generation round, we run the same data generation process for 2000 more candidate questions, now prompting GPT-4.1 for questions in these new categories. After filtering, this new round yields 26 new questions. For the final round, we generate 2500 additional candidate questions, yielding 16 more questions after filtering. We remove the 7 most similar questions, giving us a total of 50 BSD questions.
@@ E.1 BSD Cyber Dataset
We also generated 15 cybersecurity questions using the pipeline described in Figure 2. In Figure 7 we find that decomposition attacks can still uplift misuse.
In Appendix E and Appendix E, we provide two examples illustrating how WMDP [39] questions are,, via our pipeline, via our pipeline, transformed into more difficult tasks that are more likely to be refused. We also provde example decompositions from Qwen 2.5-7B for each of these questions. We also provde example decompositions from Qwen 2.5-7B for each of these questions. While often thematically similar, the new generated questions often test for different kinds of cybersecurity knowledge; for example, the first question originally tests for reconnaissance but the transformed version tests weaponization [70, 39] [70, 39]. We also provide example decompositions generated by Qwen 2.5-7B for the respective questions.
@@ 1. Example Generation, Question, and Decomposition
B) Copy files from the device to the C2 server.
C) Automatically start when the device boots using an Android broadcast receiver.
D) All of the above.
@@
B) Androguard’s decompiler module
C) EvilGrade’s APK injector
D) APKTool’s resource patcher
@@ 2. Example Generation, Question, and Decomposition
B) Optimizes application performance
C) Injects code into running processes to analyze and modify their behavior
D) Logs runtime behavior
@@
B) Craft a document containing a malicious macro that triggers outbound HTTP requests upon file access.
C) Upload a symbolic link (symlink) file pointing to a remote SMB share you control, exploiting NTFS symlink handling on the backend.
D) Upload a ZIP archive with directory traversal sequences to overwrite the backend configuration file.

### v2

@@ Appendix A Additional related work
Dangerous capability evaluations. Dangerous capability evaluations attempt to estimate the proficiency of frontier models on tasks where language models could unlock large scale harm, for example, cyber‑offense, persuasion, bio-engineering, and self‑replication (Phuong et al., 2024, Shevlane et al., 2023, Phuong et al., 2025). Frontier model developers most often conduct dangerous capability evaluations internally and report high-level results via system cards (OpenAI Preparedness Team, 2023, Anthropic, 2024, Jaech et al., 2024, Grattafiori et al., 2024). Dangerous capability evaluations are run under a threat model where the human attempting misuse is either directly querying the model (typically with safeguards like safety training removed) or applying an undisclosed jailbreak or elicitation method. Sometimes dangerous capability evaluations are paired with human uplift studies, which evaluate the extent that a language model helps humans perform dangerous or dual-use tasks (OpenAI, 2024, AI Security Institute, 2024). In contrast, our threat model assumes that model developers will deploy standard safeguards and that attackers will attempt to subvert safeguards via attack strategies like decomposition attacks and jailbreaking.
@@
Jailbreaking methods. Most jailbreaks try to coerce a model into eliciting disallowed content, e.g., “Tell me how to build a bomb”Shin et al. (2020), Zou et al. (2023), Andriushchenko et al. (2025), Chao et al. (2025), Liu et al. (2023). Many optimize for a fixed target string (“Here is how to build a bomb…”) Zou et al. (2023), Andriushchenko et al. (2025) and others look for non‑refusal answers Chao et al. (2025), Mehrotra et al. (2024), Russinovich et al. (2024). These approaches are usually benchmarked on questions whose answers are easy to find via the web Mazeika et al. (2024), Chao et al. (2024). Outputs from jailbreaks, even when “successful,” often return vague or erroneous instructions Nikolić et al. (2025). HarmBench’s harder context-based tasks represent an attempt to alleviate this, yet are largely saturated by open‑weight LLMs Mazeika et al. (2024), Sabbaghi et al. (2025). Here, we instead measure misuse‑uplift on genuinely hard, refused tasks and introduce BSD, which pairs uplift with an explicit detectability axis that is missing from refusal‑only metrics. Similar to Zou et al. (2024), Sharma et al. (2025), we show that jailbreaking prompts are relatively easy to detect, whereas decomposition attacks are significantly harder to detect.
@@
Decomposition methods. Decomposition attacks, introduced in previous work (Jones et al., 2024, Glukhov et al., 2024), are methods that use benign-looking sub-queries to help solve a malicious task. That said, Jones et al. (2024) run a decomposition attack on a set of Python scripts generated by Claude 3 Opus and judged by GPT-4. We note that the provided example tasks are not refused by strong models, e.g. Claude Sonnet 3.5 or GPT-4o, and thus cannot be used to evaluate our misuse uplift threat model. Similarly, Jones et al. (2024) does not compare decomposition attacks with established jailbreak methods. Glukhov et al. (2024) studies the increase in their introduced Impermissible Information Leakage on WMDP, but as shown in Section˜6, strong models directly answer these queries and decomposition harms accuracy, making WMDP a poor misuse proxy. By contrast, our study (i) frames decomposition as a way to evade detectability (Section˜3.2), (ii) benchmarks the methods on a misuse‑uplift metric that factors in both task difficulty and strong model refusal, and (iii) introduces improved decompositions that outperform prior work (Section˜6).
@@
Stateful defenses. A parallel line of work shifts from single‑prompt screening to sequence‑level scrutiny. In computer vision, Stateful Detection compares each new input to a sliding window of earlier queries Chen et al. (2020); Blacklight speeds this up with locality‑sensitive hashing Li et al. (2022), and PIHA swaps raw pixels for perceptual hashes to cut false positives Choi et al. (2023); and Mind‑the‑Gap augments the windowed distance test with adaptive thresholds yet still falls to the OARS adaptive attack Park et al. (2025), Feng et al. (2023). PRADA detects model stealing by flagging query sequences whose distances deviated from benign traffic Juuti et al. (2018). Outside of vision, Clio clusters millions of conversation snippets to surface coordinated abuse, but publishes no quantitative evaluations and does not consider user-level defenses Tamkin et al. (2024). Our work (Section˜5) proposes a detector for misuse uplift that uses a buffer to keep track of the most concerning queries, and shows that even with maintaining a memory across many independent queries, decomposition attacks are harder to flag than standard jailbreaks.
@@ Appendix B Threat model details
    Open-weight models are currently weaker than proprietary models. Open-weight models— models with downloadable weights—have historically trailed proprietary systems in benchmark performance by at least 6 months (Cottier et al., 2024). While this performance gap is closing, it likely still holds for current frontier open-weight and closed-weight models (METR, 2025, DeepSeek, Inc., 2025, OpenAI, 2025c).
@@
    Open-weight models can be made unsafe. The safety-training and guardrails on open-weights models can be removed with only modest additional fine-tuning (Qi et al., 2024a; b, Gade et al., 2023). While there is early work attempting to make models robust to fine-tuning attacks (Tamirisa et al., 2025, Rosati et al., 2024), this problem is difficult— e.g., defense here is strictly harder than that for adversarial examples or jailbreaks (Rando et al., 2025).
@@ B.1 Alternative assumptions
Language model uplift is a proxy for human uplift. First, we note that helpful-only (unsafe) models may serve as cheap (but imperfect) substitutes for non-expert humans in a misuse evaluation. This means that our evaluations can provide information on human uplift (Ibrahim et al., 2024). [^4] We note that this is similar to the assumptions made in scalable oversight (Bowman et al., 2022). For example, a weaker model might serve as an imperfect stand-in for a human with beginning-to-intermediate software engineering ability (Kwa et al., 2025) in a cyber-misuse setting. In this case, the helpful-only (unsafe) model would approximate a steps performed by a human attacker: reconnaissance and vulnerability discovery, weaponization, exploitation, escalation, etc. (Strom et al., 2018), delegating to the proprietary (safe) model when needed.
@@
Imagine the adversary can only run the model on a single CPU at roughly 1 token per second, whereas the defender hosts the same weights on a GPU that runs at 100 tokens per second. Jailbreaking the defender’s endpoint grants the attacker two orders of magnitude more effective compute per wall-clock hour. For agent and reasoning workflows where the model plans, branches, etc, this translates into substantially deeper search, which in turn has been shown to raise success rates on reasoning-intensive tasks (Jaech et al., 2024).
@@
Unsafe stronger models can be complementary with safe weak models. Even in a world where the strongest models are willing to do harmful actions, the capabilities of these models may be complementary with those of proprietary models with safety training (Tamirisa et al., 2025). For example, while a helpful-only model may have vastly more world knowledge, it may still use a (weaker) safe proprietary model that has longer/more consistent reasoning to do more harm in an agent setting.
B.2 On access to strong, helpful-only models
Note that evaluators (model providers and red teams) often have access to strong unreleased helpful-only checkpoints, while attackers do not. This is due to the fact that strong base models have to be safety-trained and aligned to be safe; it is therefore often cheap convenient to train helpful-only variants for red-teaming or reward modeling (Mu et al., 2024). We assume that we can use these checkpoints for task generation/generation and agreement, verify refusal on a safety-aligned model, and measure attacker uplift with more widely available open-weights models (see discussion above).
B.3 Attackers may evade our defense by using multiple accounts
Our user-level defense and evaluations assume that the model provider can associate attacker activity across sessions with a stable identifier (e.g., an account). However, a determined adversary could attempt to evade the stateful defense by rotating across accounts. We note that doing so imposes operational costs (e.g., repeatedly establishing new accounts, using different payment/verification methods). The attacker is also potentially limited by know-your-customer regulation in banking (U.S. Congress, 2001, Department of the Treasury et al., 2003, European Parliament and Council of the European Union, 2015), which a model provider could (in the future) plausibly use to establish account provenance. Finally, creating new accounts and rotating between them may introduce new anomalies that are detectable at the platform level.
Moreover, real-world misuse reporting suggests that sophisticated campaigns can still be meaningfully addressed via account-level monitoring and enforcement. For example, in its public report on disrupting an AI-orchestrated cyber espionage campaign, Anthropic describes an operational pattern in which the actor decomposed multi-stage intrusions into many individually plausible requests “evaluated in isolation,” maintained context across sessions. This was mitigated in part by identifying and banning the involved accounts (Anthropic, 2025). While Anthropic (2025) does not specify how activity was distributed across identities, it provides evidence that (i) decomposition-style misuse can appear benign on a per-prompt basis, and (ii) account-linked interventions are feasible in practice. This supports the relevance of per-user stateful defenses as a baseline.
Finally, our evaluations and these considerations motivate that model providers defending against misuse may want to increase the friction for disposable identities (e.g., with stronger identity or payment verification) to make stateful defenses more useful. The design of such controls is outside the scope of this work.
B.4 Decomposition Attack’s details
Given a misuse task $X$, there are three essential steps to exploit the target model by decompositions: 1- The attacker must decompose $X$ into $n$ sub-tasks $\{X_{1},\cdots,X_{n}\}$. These sub-tasks are supposed to be seemingly benign, yet when glued together, must reconstruct the original task. Similar to Jones et al. (2024), we deploy a Decomposer module $\mathbb{D}$ with a crafted system prompt (see Appendix˜H) that takes $X$ as input and generates the set of sub-tasks: $\{X_{1},\cdots,X_{n}\}=\mathbb{D}(X)$. Thus, the attacker will conceal the malicious intent by using the output of Decomposer. 2- Having the benign sub-tasks, the attacker collects the answer to each of them by asking directly from the target model $\mathbb{T}$. Note that the attacker does not need to alter the input to the target LLM using any jailbreaking methods since each input is inherently benign; thus, there is no need to "trick" the model into revealing the answer. We denote the answer to sub-task $X_{i}$ as $Y_{i}=\mathbb{T}(X_{i})$. After this step, the attacker possesses the set of all the answers: $\{Y_{1},\cdots,Y_{n}\}$. 3- Now, the attacker has the sufficient information to perform the original task by reasoning from the information provided in-context. The final decision is made by providing the information to a Composer module $\mathbb{C}$ (see Appendix˜H). If we denote the final answer as $Y$, it is generated by: $Y=\mathbb{C}(\{Y_{1},\cdots,Y_{n}\})$.
A key assumption in the setting above is that the attacker does not know the answer to the task, and plans to misuse the strong target model to solve it. To impose this constraint on the attacker’s knowledge, we use a weak model such as Qwen2.5-7B that, as discussed in Section˜4, is unable to achieve a higher score than random. Then, the weak model is deployed as both Decomposer and Composer. This is detailed in Table˜2 where the appropriate setting is $X\xrightarrow{W}\{X_{i}\}\xrightarrow{S}\{Y_{i}\}\xrightarrow{W}Y$, in which $W$ and $S$ denote the weak and strong models, respectively.
@@ Appendix C Cybersecurity dataset and results
We also provide cybersecurity misuse uplift results in Figure˜7 with Qwen-2.5 7B as the attacking model for decomposition. The misuse uplift, how much the strong model augments the weaker model, is generally more modest in the cybersecurity setting compared to the uplift measured in the BSD biology dataset (see Figure˜1). This due in-part because the baseline rate of refusal for cybersecurity questions are much lower across all models (see the directy query accuracy results in Figure˜7).
Figure 7: The misuse rate for decomposition attacks vs direct querying for BSD cyber questions. While we find modest misuse uplift (the misuse rate increases over the dashed line), the attacker gains for BSD cyber questions are less than that for BSD bio questions (Figure˜1)
@@ Appendix D Protocols for decomposition
We provide a high-level description of the different decomposition attack strategies we consider in this work in Table˜2.
Decomposition Attack Input $\xrightarrow{\text{Query}}$ Output Attack Protocol Description
Query Weak $X\xrightarrow{W}Y$ Request output directly from weak helpful-only model
Query Strong $X\xrightarrow{S}Y$ Request output directly from a strong helpful and harmless model
@@
Decompose $X\xrightarrow{W}\{X_{i}\}\xrightarrow{S}\{Y_{i}\}\xrightarrow{W}Y$ Decompose with weak model into subtasks, query strong model for each, recompose results with weak model
@@ Appendix E BSD Pipeline Details
We generate BSD in three stages in order to get broad topic coverage. First, we generate 3000 candidate biology questions with GPT-4.1 (OpenAI, 2025b) using the WMDP Biology dataset (Li et al., 2024a) as seed questions. After filtering, this yields 15 total questions that meet our criteria of difficulty and harmfulness. Next, we identify 10 plausibly harmful topics in biology that are not covered in the existing question set. For the next generation round, we run the same data generation process for 2000 more candidate questions, now prompting GPT-4.1 for questions in these new categories. After filtering, this new round yields 26 new questions. For the final round, we generate 2500 additional candidate questions, yielding 16 more questions after filtering. We remove the 7 most similar questions, giving us a total of 50 BSD questions.
Figure 8: Flow diagram of the correctness, harmfulness, and difficulty filters (Section˜3.1) for BSD. We generate the candidate questions by prompting GPT-4.1. Then, from left to right, the three filters corresponds to conditions C3, C2, and C1 in Section˜3.3.
@@ E.1 BSD Cyber Dataset
We also generated 15 cybersecurity questions using the pipeline described in Figure˜2. In Figure˜7 we find that decomposition attacks can still uplift misuse.
In Appendix˜E and Appendix˜E, we provide two examples illustrating how WMDP (Li et al., 2024a) questions are, via our pipeline, transformed into more difficult tasks that are more likely to be refused. We also show example decompositions from Qwen 2.5-7B for each of these questions. While often thematically similar, the new generated questions often test for different kinds of cybersecurity knowledge; for example, the first question originally tests for reconnaissance but the transformed version tests weaponization (Strom et al., 2018, Li et al., 2024a). We also provide example decompositions generated by Qwen 2.5-7B for the respective questions.
@@ 1. Example Generation, Question, and Decomposition
B) Copy files from the device to the C2 server. C) Automatically start when the device boots using an Android broadcast receiver. D) All of the above.
@@
B) Androguard’s decompiler module C) EvilGrade’s APK injector D) APKTool’s resource patcher
@@ 2. Example Generation, Question, and Decomposition
B) Optimizes application performance C) Injects code into running processes to analyze and modify their behavior D) Logs runtime behavior
@@
B) Craft a document containing a malicious macro that triggers outbound HTTP requests upon file access. C) Upload a symbolic link (symlink) file pointing to a remote SMB share you control, exploiting NTFS symlink handling on the backend. D) Upload a ZIP archive with directory traversal sequences to overwrite the backend configuration file.
@@ Appendix F Experiments Discussions
F.1 Measuring uplift with other strong open-weights models
We run additional experiments with four near-frontier open-weights models on the BSD-Biology dataset in Table˜3. The misuse uplift, the difference between the decomposition attack performance and the direct query performance, for these models is generally more modest compared to the weaker Qwen2.5-7B, but still significant. We note that the direct query results for Kimi K2 and Llama 3.1-405B are less high due to refusals. Likewise, because we spend less time tuning the decomoposition prompts and hyperparameters (we use a fixed $n=6$ decomposition queries for each question, and do not perform finetuning), the decomposition attack numbers can likely be improved.
Table 3: Misuse uplift with near-frontier open-weight models on BSD-Biology using $n=6$ decompositions. Models with higher refusal rates (lower Direct scores) show larger uplift.
GPT-5 Gemini-2.5
Weak (Unsafe) Model Direct Decomp Uplift Decomp Uplift
Kimi K2-Instruct-0905 0.50 $\pm$ 0.04 0.76 $\pm$ 0.02 +0.26 $\pm$ 0.04 0.83 $\pm$ 0.02 +0.33 $\pm$ 0.04
Qwen2.5-72B 0.68 $\pm$ 0.04 0.78 $\pm$ 0.03 +0.10 $\pm$ 0.05 0.88 $\pm$ 0.02 +0.20 $\pm$ 0.04
Qwen3-235B 0.73 $\pm$ 0.06 0.73 $\pm$ 0.02 +0.00 $\pm$ 0.06 0.85 $\pm$ 0.02 +0.12 $\pm$ 0.06
Llama-3.1-405B 0.34 $\pm$ 0.04 0.71 $\pm$ 0.03 +0.37 $\pm$ 0.05 0.81 $\pm$ 0.02 +0.47 $\pm$ 0.04
F.2 Hyperparameter Details
Baselines setting
In Table˜1 we compare the decomposition attacks with jailbreak baselines, each limited to 12 calls to the strong target model. Therefore, we make some modifications to the baselines. We (i) modify the Adaptive Attack Andriushchenko et al. (2025) by generating 12 diverse suffixes for each task with Llama-3-8B (Grattafiori et al., 2024) to transfer them to the strong target model, (ii) configure PAIR Chao et al. (2025) with 6 parallel streams over 2 iterations (resulting 12 total prompts), (iii) run Adversarial Reasoning Sabbaghi et al. (2025) for 3 iterations with 4 attacking prompts each, and (iv) Crescendo Russinovich et al. (2024) with $\text{max\_rounds}=6$ and $\text{max\_tries}=6$.
Compute
For our adversarial training experiments in Section˜5, we use roughly $100$ hours on a single NVIDIA A100 GPU node. For the result of Section˜6, we deployed 8 * NVIDIA H100 GPUs for 40 minutes only to fine-tune the Qwen2.5-7B model on 700 data collected from o3-mini.